### v3

@@ Abstract
Electronic Health Records (EHRs) offer considerable potential for clinical prediction, but their complexity and heterogeneity present significant challenges for traditional machine learning methods. Recently, domain-specific EHR foundation models trained on large volumes of unlabeled EHR data have shown improved predictive accuracy and generalization. However, their development is constrained by limited access to diverse, high-quality datasets, and by inconsistencies in coding standards and clinical practices. In this study, we explore the use of general-purpose Large Language Models (LLMs) to encode EHR into high-dimensional representations for downstream clinical prediction tasks. We convert structured EHR data into markdown-formatted plain text documents by replacing medical codes with natural language descriptions. This enables the use of LLMs and their extensive semantic understanding and generalization capabilities as effective encoders of EHRs without requiring access to private medical training data. We show that LLM-based embeddings can often match or even surpass the performance of a specialized EHR foundation model, CLMBR-T-Base, across 15 diverse clinical tasks from the EHRSHOT benchmark. To demonstrate generalizability, we further evaluate the approach on the UK Biobank (UKB) cohort, a population distinct from that used to train CLMBR-T-Base. Notably, one of the tested LLM-based models achieves superior performance for disease onset, hospitalization, and mortality prediction, highlighting robustness to shifts in patient populations. Our findings suggest that repurposed general-purpose LLMs for EHR encoding provide a scalable and generalizable alternative to domain-specific models for clinical prediction.
@@
However, machine learning on EHR data poses significant challenges due to its inherent complexity. EHR data is characterized by variable-length sequences of patient visits, irregular sampling intervals, missing entries, heterogeneous and noisy information, and a wide range of hierarchical medical concepts [11]. As a result, deep learning models often achieve only modest improvements over traditional methods such as logistic regression or tree-based methods [6, 12, 13]. To mitigate these issues, recent approaches have employed large-scale foundation models that are pre-trained on unlabeled EHR data using unsupervised learning [14]. Many of these models adopt strategies from natural language processing, such as masked-word prediction as in BERT [15] or autoregressive next-word prediction as in GPT [16]. Treating EHR data as sequences of medical codes enables analogous methods such as masked code prediction [17, 18, 12, 19] or next code prediction [13]. However, these techniques also face significant limitations: coding standards and healthcare practices differ strongly across sites, and interoperable EHR foundation models would likely need to be trained on a wide variety of EHR datasets, which is difficult to achieve due to the sensitivity of healthcare data. Therefore, the development of EHR-specific foundation models remains constrained by the limited size and restricted availability of EHR data.
@@
In contrast, LLMs benefit from training on vast general-purpose text corpora and a broad range of natural language tasks [20, 21]. This extensive pre-training enables their language comprehension and allows them to capture domain-agnostic patterns that can be adapted for healthcare applications. Consequently, LLMs have demonstrated strong performance in extracting medical concepts [22], summarizing medical texts [23], and predicting medical outcomes [24] even in low-resource settings. However, most modern LLMs, such as GPT [25] or Llama [26], use a decoder-only transformer architecture, which complicates the generation of robust text representations. To overcome this limitation, recent work has introduced methods to convert decoder-only LLMs into effective embedding models for downstream prediction tasks [27, 28, 29, 30]. Additionally, these state-of-the-art models offer an increased context window, making them well-suited for handling long inputs such as serialized EHR data.
@@
In this study, we evaluate whether general-purpose LLM embedding models can effectively encode EHR data for clinical prediction [31] (see Fig. 1). To this end, we convert structured EHR records into plain text summaries that highlight key patient information at prediction time. Using two state-of-the-art LLM embedding models, GTE-Qwen2-7B-Instruct (GTE-Qwen2-7B) [32, 30] and LLM2Vec-Llama-3.1-8B-Instruct (LLM2Vec-Llama-3.1-8B) [27, 33], we generate high-dimensional EHR embeddings that serve as inputs to logistic regression classifiers across 15 clinical tasks from the EHRSHOT benchmark. We focus on the performance in few-shot settings to evaluate the generalization ability of this approach and conduct extensive ablation studies to identify the specific components that drive the LLM ’s effectiveness. Finally, we validate our approach on the UKB for predicting mortality, hospitalization and onset of 23 diseases [34] to assess its robustness and compare the generalizability with EHR-specific foundation models.
@@ 2.1 Experimental Setup
Our primary analysis was conducted using the EHRSHOT database, which contains EHRs from 6,739 adult patients treated at Stanford Health Care and Lucile Packard Children’s Hospital between 1990 and 2023. This dataset includes 921,499 visits and over 41.6 million clinical events [31], and serves as a standardized benchmark with 15 clinical prediction tasks organized into 4 task groups, along with predefined splits and publicly available code. Table 1 summarizes cohort statistics, and task details are shown in Table 2. Further information on task definitions and preprocessing is provided in Section 4.1.
@@
For external validation, we used the UKB, a population-based cohort comprising 502,489 UK participants [35, 36]. This allowed us to evaluate the generalizability of our approach across healthcare systems, especially since the CLMBR-T-Base model was trained on data from the same hospital system as EHRSHOT. Due to structural differences between the datasets, not all EHRSHOT tasks were transferable to the UKB (see Section 4.8). We focused on predicting one-year risk of hospitalization, mortality, and the onset of 23 chronic diseases [34]. The processed UKB subset used in our study consisted of 387,464 patients, approximately 19.5 million visits, and over 72 million clinical events. Table 6 provides task descriptions and label distributions.
@@
To prepare EHR data for use with LLMs, we serialized structured patient records into plain text formatted in Markdown, with a maximum length of 4,096 tokens, prioritizing more recent data. For EHRSHOT, serialization began with a reference prediction date to which all event dates were normalized, followed by demographics, recent values for 24 key LOINC concepts across Body Metrics, Vital Signs, and Lab Results, including units and classification (low/normal/high). These were followed by a summary of patient visits and reverse-chronological listings of visit-level events such as conditions, medications, and procedures. An example EHR text serialization is shown in Fig. 2, and a detailed explanation appears in Section 4.2. For the UKB, a similar serialization approach was used, with event dates normalized to each patient’s recruitment date and exclusion of detailed measurements due to data limitations.
@@
We evaluated two instruction-tuned LLM embedding models, GTE-Qwen2-7B [32, 30] and LLM2Vec-Llama-3.1-8B [33, 27], using task-specific prompts (e.g., “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”, see Table 8). Our primary objective was to assess how general-purpose LLM-based embedding models, trained on public text data, perform in encoding EHRs for clinical prediction tasks. For comparison, we included CLMBR-T-Base, a 141-million-parameter autoregressive foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine [13, 31]. For each model, we computed patient-level embeddings and trained a logistic regression classifier on the training split, with hyperparameters selected using a validation set. As a baseline, we also included a Gradient Boosted Machine (GBM) classifier trained on count-based features, which has previously demonstrated strong performance on EHR tasks [31]. See Section 4.4 for additional details on model architectures.
@@
| LLM-Embedding Models | |||||
| GTE-Qwen2-7B | |||||
| GTE-Qwen2-1.5B | |||||
| LLM2Vec-Llama-3.1-8B | |||||
| LLM2Vec-Llama 2 1.3B | |||||
| LLM-Embedding Model + EHR Foundation Model | |||||
| GTE-Qwen2-7B + CLMBR-T-Base | |||||
| LLM2Vec-Llama-3.1-8B + CLMBR-T-Base | |||||
@@
| Counts-based + GBM | |||||
| Encoder Language Models | |||||
| DeBERTaV3 large | |||||
| DeBERTa V3 base | |||||
| BERT large | |||||
| BERT base | |||||
@@
| Counts-based + GBM | ||||
Using all available training and validation examples, the GTE-Qwen2-7B model outperformed CLMBR-T-Base in two of the four task categories: lab test result prediction and assignment of new diagnoses (see Table 3). In contrast, CLMBR-T-Base achieved higher performance on tasks related to operational outcomes and chest X-ray findings. LLM2Vec-Llama-3.1-8B embeddings showed slightly lower performance than GTE-Qwen2-7B, exceeding CLMBR-T-Base only in the assignment of new diagnoses task group. Both LLM-based models were often on par with or outperformed the GBM-based baseline, which is considered a strong baseline for EHR prediction tasks [12, 13]. To assess whether the models encode orthogonal information, we concatenated CLMBR-T-Base embeddings with LLM-based embeddings and observed substantial performance improvements for both LLM models. The combined embeddings achieved an average area under the receiver operating characteristic curve (AUROC) of 0.786 (0.762–0.811) for GTE-Qwen2-7B, up from 0.755 (0.730–0.780), and 0.773 (0.747–0.799) for LLM2Vec-Llama-3.1-8B, up from 0.734 (0.707–0.762). The smaller GTE-Qwen2-1.5B model achieved an average performance of 0.748 (0.722–0.774), only marginally lower than its larger counterpart, suggesting it is a viable lightweight alternative for EHR embeddings. In contrast, LLM2Vec-Llama-2-1.3B performed notably worse than the larger version, potentially due to limitations of the older Llama 2 architecture. Figure 3 shows a general trend that larger models tend to yield better performance across tasks. Among the evaluated models, CLMBR-T-Base remained the most parameter-efficient relative to predictive performance. Nonetheless, it is important to note that the LLM-based embedding models were not specifically optimized for structured EHR data, which may explain their performance differences.
On the UKB dataset, GTE-Qwen2-7B consistently outperformed CLMBR-T-Base across all three prediction task groups (see Table 4). LLM2Vec-Llama-3.1-8B surpassed CLMBR-T-Base only in the task group related to the assignment of new diagnoses. The performance drop of CLMBR-T-Base was anticipated, as the UKB represents an out-of-domain dataset with numerous disease codes not included in the model’s original vocabulary. Although we applied a careful mapping of UKB Observational Medical Outcomes Partnership (OMOP) concepts to those supported by CLMBR-T-Base to improve compatibility and predictive performance, many concepts remained unmapped due to CLMBR-T-Base’s limited vocabulary. This limitation underscores a key challenge in adapting domain-specific models to external datasets, even with considerable manual effort. In contrast, the LLM-based models required only a mapping from clinical concepts to their natural language descriptions, which were directly incorporated into the Markdown-based serialization. This simplified integration process substantially reduced implementation overhead and emphasized the strong generalization ability of LLM-based encodings across diverse coding systems and healthcare domains.
@@ 2.3 LLM Embeddings Achieve Strong Performance in Low-Data Regimes
To evaluate model performance with limited training data, we conducted experiments in a few-shot setting using small numbers of training examples. The experiments followed the EHRSHOT task definitions [31]. Fig. 5 illustrates the aggregated AUROC across all subtasks within the four task categories for varying numbers of training examples. LLM embedding models performed strongly in the few-shot setting, suggesting effective transfer from general text pretraining to serialized EHR data. Notably, GTE-Qwen2-7B consistently outperformed CLMBR-T-Base across all training sizes for lab test result prediction and assignment of new diagnoses. The strong results for lab test predictions indicate that the inclusion of the last three lab values in the EHR serialization provides highly predictive features, which GTE-Qwen2-7B effectively encodes. In contrast, LLM2Vec-Llama-3.1-8B showed comparatively weaker performance in the few-shot regime. It only surpassed CLMBR-T-Base in a limited subset of tasks, primarily those involving the assignment of new diagnoses. Nonetheless, both LLM-based models consistently outperformed the count-based baseline across most configurations, underscoring the benefits of extensive pretraining. Consistent with previous findings [31], the largest performance improvements over the baseline were observed at intermediate training set sizes. As the number of labeled data samples increased, the relative advantage of pre-trained LLM-based models diminished, highlighting the reduced marginal gains of foundation models in high-data regimes. The area under the precision-recall curve (AUPRC) results are shown in Fig. 7, with task-specific AUROC and AUPRC results provided in Fig. 8 and Fig. 9 in the Appendix.
@@
In contrast to EHRSHOT, the LLM-based methods outperformed CLMBR-T-Base in most prediction tasks on the UKB dataset. For the assignment of new diagnoses, which includes 23 tasks, both GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B outperformed CLMBR-T-Base across all sample sizes (see Fig. 10 in the Appendix). For hospitalization and mortality prediction, CLMBR-T-Base showed stronger performance in the very low-data regime. However, with more samples the performance of the LLM embedding models increased and GTE-Qwen2-7B outperformed CLMBR-T-Base. Figure 5 presents the mean AUROC across all task groups for both EHRSHOT and UKB, highlighting the observed differences. On EHRSHOT, GTE-Qwen2-7B generally matched CLMBR-T-Base in performance, while LLM2Vec-Llama-3.1-8B fell between these models and the count-based baseline. On the UKB dataset, LLM-based models and CLMBR-T-Base performed comparably with very limited training data of up to four examples. With additional training data, GTE-Qwen2-7B surpassed both CLMBR-T-Base and LLM2Vec-Llama-3.1-8B, while the latter closely approached the performance of CLMBR-T-Base. These findings underscore the capability of general-purpose LLM embeddings to rival or exceed the performance of domain-specific models, particularly in settings characterized by domain shift or limited labeled data. Detailed task-specific results are available in Fig. 12 and Fig. 13 in the Appendix.

### v4

@@ Abstract
Electronic Health Records (EHRs) offer considerable potential for clinical prediction, but their complexity and heterogeneity present significant challenges for traditional machine learning methods. Recently, domain-specific EHR foundation models trained on large volumes of unlabeled EHR data have shown improved predictive accuracy and generalization. However, their development is constrained by limited access to diverse, high-quality datasets, and inconsistencies in coding standards and clinical practices. In this study, we explore the use of general-purpose Large Language Models (LLMs) to encode EHR into high-dimensional representations for downstream clinical prediction tasks. We convert structured EHR data into Markdown-formatted plain-text documents by replacing medical codes with natural language descriptions. This enables the use of LLMs and their extensive semantic understanding and generalization capabilities as effective encoders of EHRs without requiring access to private medical training data. We show that LLM-based embeddings can often match or even surpass the performance of a specialized EHR foundation model, CLMBR-T-Base, across 15 diverse clinical tasks from the EHRSHOT benchmark. Critically, our approach requires no institution-specific training and can incorporate any medical code with a text description, whereas existing EHR foundation models operate on fixed vocabularies and can only process codes seen during pretraining. To demonstrate generalizability, we further evaluate the approach on the UK Biobank (UKB) cohort, out-of-domain for CLMBR-T-Base, whose fixed vocabulary covers only 16% of UKB codes. Notably, an LLM-based model achieves superior performance for prediction of disease onset, hospitalization, and mortality, indicating robustness to population and coding shifts. Our findings suggest that repurposed general-purpose LLMs for EHR encoding provide a scalable and flexible alternative to domain-specific models for clinical prediction.
@@
However, machine learning on EHR data poses significant challenges due to its inherent complexity. EHR data is characterized by variable-length sequences of patient visits, irregular sampling intervals, missing entries, heterogeneous and noisy information, and a wide range of hierarchical medical concepts [11]. As a result, deep learning models often achieve only modest improvements over traditional methods such as logistic regression or tree-based models for EHR prediction tasks [6, 12, 13]. To mitigate these issues, recent approaches have employed large-scale foundation models that are pretrained on unlabeled EHR data using unsupervised learning [14]. Many of these models adopt strategies from natural language processing, such as masked word prediction as in BERT [15] or autoregressive next word prediction as in GPT [16]. Treating EHR data as sequences of medical codes enables analogous methods such as masked code prediction [17, 18, 12, 19] or next code prediction [13, 20, 21]. However, code-based EHR foundation models face two fundamental obstacles to interoperability and generalization: site-specific coding practices and fixed vocabularies learned during pretraining. For example, CLMBR-T-Base [13] supports only 26,249 unique codes from its training corpus, and when applied to the UKB with 50,702 unique medical codes, only 7,969 (16%) could be mapped, leaving 84% of codes unseen by the model. Achieving interoperability would require pretraining on diverse EHR datasets from many institutions, which is difficult due to the sensitivity of healthcare data. This motivates models that operate on natural-language descriptions of clinical codes, which avoid fixed vocabularies and transfer more readily across institutions.
@@
LLMs benefit from pretraining on vast general-purpose text corpora and a broad range of natural-language tasks [22, 23]. This extensive pretraining enables their language comprehension and allows them to capture domain-agnostic patterns that can be adapted for healthcare applications. Consequently, LLMs have demonstrated strong performance in extracting medical concepts [24], summarizing medical texts [25], and predicting medical outcomes [26], even in low-resource settings. Recent work extends LLMs to structured EHR by serializing records into text and either using model generations for prediction [27, 28, 29, 30, 31, 32] or extracting fixed-dimensional embeddings for downstream classifiers [33, 34, 35]. While both paradigms can be competitive with common baselines, many prior studies only use short context windows and evaluate on private or emergency-department cohorts, limiting longitudinal coverage and external validity [33, 35, 34] (see Section 4.7). Moreover, most modern LLMs, such as GPT [36] and Qwen [37, 38], use decoder-only transformer architectures trained with left-to-right objectives, which are not optimized for representation learning. To address this limitation, recent work converts decoder-only LLMs into effective LLM embedding models via contrastive learning or related techniques [39, 40, 41, 42, 43]. Additionally, these state-of-the-art models offer large context windows, making them well-suited for handling long inputs such as serialized EHR data.
@@
In this study, we evaluate whether general-purpose LLM embedding models can effectively encode EHR data for clinical prediction [44] (see Fig. 1). To this end, we convert structured EHR records into plain-text summaries that highlight key patient information at prediction time. Using a state-of-the-art LLM embedding model, Qwen3-Embedding-8B (Qwen3-Emb-8B) [38, 43] with a context size of 8,192 tokens, we generate high-dimensional EHR embeddings that serve as inputs to logistic regression classifiers across 15 clinical tasks from the EHRSHOT benchmark. We analyze the performance in the few-shot setting, where the downstream logistic regression is trained with only a small number of labeled examples, to evaluate the generalization of this approach, and we conduct extensive ablation studies to identify the components that drive the LLM ’s effectiveness. Finally, we validate our approach on the UKB for predicting mortality, hospitalization, and the onset of 23 diseases [45] to assess robustness and compare its generalizability against EHR-specific foundation models.
@@ 2.1 Experimental Setup
Our primary analysis was conducted using the EHRSHOT benchmark, which contains EHRs from 6,739 adult patients treated at Stanford Health Care and Lucile Packard Children’s Hospital between 1990 and 2023. This dataset includes 921,499 visits and over 41.6 million clinical events [44], and serves as a standardized benchmark with 15 clinical prediction tasks organized into 4 task groups, along with predefined splits and publicly available code. Table 1 summarizes cohort statistics, and task details are shown in Table 2. Further information on task definitions and preprocessing is provided in Section 4.1.
@@
For external validation, we used the UKB, a population-based cohort comprising 502,489 UK participants [46, 47]. This allowed us to evaluate the generalizability of our approach across healthcare systems, especially since the CLMBR-T-Base model was trained on data from the same hospital system as EHRSHOT. Due to structural differences between the datasets, not all EHRSHOT tasks were transferable to the UKB (see Section 4.10). We thus focused on predicting one-year risk of hospitalization, mortality, and the onset of 23 diseases [45] for the external validation in the UKB. The processed UKB subset used in our study consisted of 387,464 patients, approximately 19.5 million visits, and over 72 million clinical events. Table 7 provides task descriptions and label distributions.
@@
To prepare EHR data for use with LLMs, we serialized structured patient records into plain text formatted in Markdown, with a maximum length of 8,192 tokens, prioritizing more recent data. For EHRSHOT, serialization began with a reference prediction date to which all event dates were normalized, followed by demographics, recent values for 24 key Logical Observation Identifiers Names and Codes (LOINC) concepts across Body Metrics, Vital Signs, and Lab Results, including units and classification (low/normal/high). These were followed by a summary of patient visits and reverse-chronological listings of visit-level events such as conditions, medications, and procedures. An example EHR text serialization is shown in Fig. 2, and a detailed explanation appears in Section 4.2. For UKB, a similar serialization approach was used, with event dates normalized to each patient’s recruitment date and exclusion of detailed measurements due to data limitations. In some experiments, we used a 4,096-token limit to accommodate runtime constraints.
@@
We evaluated three instruction-tuned LLM embedding models, Qwen3-Embedding-8B (Qwen3-Emb-8B) [38, 43], GTE-Qwen2-7B-Instruct (Qwen2-Emb-7B) [37, 42], and LLM2Vec-Llama-3.1-8B-Instruct (Llama3.1-LLM2Vec-8B) [48, 39], with a primary focus on the first model due to its recency. We used task-specific prompts (e.g., “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”, see Table 9). Our primary objective was to assess how general-purpose LLM-based embedding models, trained on public text data, perform in encoding EHRs for clinical prediction tasks. For comparison, we included CLMBR-T-Base, a 141-million-parameter autoregressive foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine [13, 44]. Since the EHRSHOT benchmark was derived from the same hospital system, CLMBR-T-Base can be considered an in-domain baseline for this benchmark. Notably, the validation and test splits of the benchmark were kept completely separate from the model pretraining (see Fig. 1 in [44]). For each model, we computed patient-level embeddings and trained a logistic regression classifier on the training split, with hyperparameters selected using a validation set. Following the EHRSHOT protocol, we used the same simple logistic-regression head for all embedding-based representations to assess embedding quality without adding classifier complexity or overfitting risk in the few-shot setting. As a baseline, we also trained a Gradient Boosted Machine (GBM) classifier on counts of medical concepts aggregated over predefined intervals before prediction, a method that has previously shown strong performance on EHR tasks [44]. To test the effect of a GBM head, we additionally trained a GBM on Qwen3-Emb-8B embeddings, which yielded worse results (see Table 11). See Section 4.4 for additional details on model architectures.
Using all training and validation examples, Qwen3-Emb-8B performed nearly on par with the in-domain EHR foundation model CLMBR-T-Base on EHRSHOT, with an area under the receiver operating characteristic curve (AUROC) of 0.761 (0.733–0.788) compared to 0.769 (0.746–0.792). The LLM embedding model outperformed CLMBR-T-Base in two of the four task categories: lab test result prediction and assignment of new diagnoses (see Table 3). Qwen3-Emb-8B matched or exceeded the GBM-based counts baseline, a strong model for EHR prediction tasks [12, 13]. To assess whether the models encode complementary information, we concatenated the LLM-based and CLMBR-T-Base embeddings and observed a substantial improvement to 0.784 (0.758–0.811). We also evaluated the Multiple Embedding Model for EHR (MEME), which encodes different sections of the EHR serialization separately, yielding 0.771 (0.745–0.797) for Qwen3-Emb-8B and 0.734 (0.705–0.764) for MedBERT [34]. However, MEME required eight times as many model inferences, and the overall performance of the original MedBERT model remained considerably lower than Qwen3-Emb-8B on the full EHR serialization. Smaller model variants showed decreased average AUROC: 0.739 (0.713–0.764) for Qwen3-Emb-4B and 0.722 (0.693–0.752) for Qwen3-Emb-0.6B. The best encoder-only language models were BERT variants pretrained on medical data. Nevertheless, they performed considerably worse than the LLM embedding models, with 0.707 (0.680–0.734) for ClinicalBERT and 0.694 (0.661–0.727) for MedBERT. Results for all models are provided in Table 10. We plotted parameter counts versus average performance of all language models with different sizes in Fig. 3, suggesting that the performance of the Qwen3- and Qwen2-based embedding models scales with model size. We only observe a slight performance trend for the encoder-only model DeBERTa, while for BERT the performance declines with larger models. However, these encoder-only models were applied with chunked 512-token inputs, which may have contributed to this effect (see Section 4.4). Among the evaluated models, CLMBR-T-Base remained the most parameter-efficient relative to predictive performance. Nonetheless, the LLM-based embedding models were not specifically optimized for structured EHR data, which may explain their performance differences.
@@ 2.2 General-Purpose LLM Embeddings Rival Domain-Specific EHR Models
| Count-based + GBM | |||||
| LLM Embedding Models | |||||
| Qwen3-Emb-8B | |||||
| Qwen3-Emb-4B | |||||
| Qwen3-Emb-0.6B | |||||
| LLM Embedding Model + EHR Foundation Model [44] | |||||
| Qwen3-Emb-8B + CLMBR-T-Base | |||||
| Multiple Embedding Model for EHR (MEME) [34] with Linear Head | |||||
| Qwen3-Emb-8B MEME | |||||
| MedBERT MEME | |||||
| Encoder Language Models with Chunked Inputs | |||||
@@
| Count-based + GBM | ||||
| LLM Embedding Model | ||||
| Qwen3-Emb-8B | ||||
On the UKB dataset, Qwen3-Emb-8B consistently outperformed CLMBR-T-Base across all three prediction task groups with an overall performance of 0.748 (0.736–0.760) compared to 0.713 (0.698–0.728) (see Table 4). The performance drop of CLMBR-T-Base was expected because UKB is out-of-domain and contains many disease codes not included in the model’s pretraining vocabulary. Of UKB ’s 50,702 unique medical codes, only 7,969 (16%) mapped to CLMBR’s fixed vocabulary of 26,249 codes. Consequently, 84% of codes and their associated clinical signal were effectively invisible to CLMBR-T-Base. We applied careful Observational Medical Outcomes Partnership (OMOP) concept mapping to maximize coverage, but fundamental vocabulary constraints remained. In contrast, the LLM-based approaches can utilize all 50,702 codes by mapping each to a natural-language description during serialization. This ability to incorporate any code with a semantic description, without retraining, is a practical advantage for cross-institution deployment where coding practices vary substantially.
@@ 2.3 LLM Embeddings Achieve Strong Performance in Low-Data Regimes
To evaluate model performance with limited training data, we conducted experiments in a few-shot setting using small numbers of training examples. The experiments followed the EHRSHOT task definitions [44]. Fig. 5 illustrates the aggregated AUROC across all subtasks within the four task categories for varying numbers of training examples. The LLM embedding model performed strongly in the few-shot setting, suggesting effective transfer from general text pretraining to serialized EHR data. Notably, Qwen3-Emb-8B consistently outperformed CLMBR-T-Base across all training sizes for lab-test result prediction and assignment of new diagnoses. The strong results for lab-test predictions indicate that inclusion of the last three lab values in the EHR serialization provided highly predictive features, which Qwen3-Emb-8B effectively encoded. Additionally, Qwen3-Emb-8B consistently outperformed the count-based baseline across most configurations, underscoring the benefits of extensive pretraining. Consistent with previous findings [44], the largest performance improvements over the baseline were observed at intermediate training set sizes. As the number of labeled data samples increased, the relative advantage of pretrained LLM-based models diminished, highlighting the reduced marginal gains of foundation models in high-data regimes. The area under the precision-recall curve (AUPRC) results are shown in Fig. 9, with task-specific AUROC and AUPRC results provided in Fig. 10 and Fig. 11 in the Appendix.
@@
In contrast to EHRSHOT, the LLM-based method outperformed CLMBR-T-Base in most prediction tasks on the UKB dataset. For the assignment of new diagnoses, which includes 23 tasks, Qwen3-Emb-8B outperformed CLMBR-T-Base across all sample sizes larger than two (see Fig. 12 in the Appendix). For mortality prediction and operational outcomes, CLMBR-T-Base showed stronger performance in the very low-data regime of up to four training samples. However, with more samples the performance of the LLM embedding models increased, and Qwen3-Emb-8B outperformed CLMBR-T-Base. Figure 5 presents the AUROC across all task groups for both EHRSHOT and UKB, highlighting the observed differences. On EHRSHOT, Qwen3-Emb-8B generally matched CLMBR-T-Base in performance and consistently outperformed the count-based baseline. On the UKB dataset, the LLM-based model and CLMBR-T-Base performed comparably with very limited training data of up to four examples. With additional training data, Qwen3-Emb-8B consistently outperformed CLMBR-T-Base. These findings underscore the capability of general-purpose LLM embeddings to rival or exceed the performance of domain-specific models, particularly in settings characterized by domain shift or limited labeled data. Detailed task-specific results are available in Fig. 14 and Fig. 15 in the Appendix.
2.4 Effect of the EHR Serialization on LLM-Based Embedding Performance